Jump to main content.


Drake Chemical

EPA Daily Situation Reports for October 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Saturday, October 31, 1998

FOR DATES: From 7:00 a.m. Friday, Oct. 30, 1998 to 7:00 a.m. Saturday, Oct. 31, 1998.

WEATHER: Clear. The temperature range was 300F to 600F.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 956.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Gregory A. Beauseigneur for David Modricker (USACE)


Friday, October 30, 1998

FOR DATES: From 7:00 a.m. Thursday, Oct. 29, 1998 to 7:00 a.m. Friday, Oct. 30, 1998.

WEATHER: Clear and cool. The temperature range was 30oF to 59oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 947.0 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to a flameout of a kiln burner and a Secondary Combustion Chamber burner. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical report (bot-ash-10-25-98-260) for bin 1. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-10-22-98-048) for bin B. Test results indicated that the fly ash materials met the project requirements for restricted backfilling (i.e. above elevation 553) due to the concentrations of cadmium, chromium, and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: A site tour was given to a group of students from East Lycoming High School.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, October 29, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 28, 1998 to 7:00 a.m. Thursday, Oct. 29, 1998.

WEATHER: Overcast with period of light rain. The temperature range was 38oF to 67oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 957.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation was performed at the southeast corner of the exclusion zone, adjacent to the sheetpiling. Samples of ground water from the exclusion zone excavation bottom were collected. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-10-24-98-259) for bin 2. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (ie: above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: A site tour was given to a small group of local residents.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, October 28, 1998

FOR DATES: From 7:00 a.m. Tuesday, Oct. 27, 1998 to 7:00 a.m. Wednesday, Oct. 28, 1998.

WEATHER: Mild with evening rain. The temperature range was 45oF to 63oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 959.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation was performed at the southeast corner of the exclusion zone, adjacent to the sheetpiling. Continued with the real time perimeter air monitoring (PAM) program and began the weekly OSHA method 93 and NIOSH 5518 testing events for BNA at the perimeter and off-site air monitoring stations. Received bottom ash analytical report (bot-ash-10-23-98-258) for bin 3. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Jared Olsen (USACE), Matt Radek (OHM), Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, October 27, 1998

FOR DATES: From 7:00 a.m. Monday, Oct. 26, 1998 to 7:00 a.m. Tuesday, Oct. 27, 1998.

WEATHER: Partly sunny with evening rain. The temperature range was 39oF to 67oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 948.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation was performed at the southeast corner of the exclusion zone, adjacent to the sheetpiling. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-10-21-98-256 and bot-ash-10-22-98-257) for bin 5 and bin 4 respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. at or above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, October 26, 1998

FOR DATES: From 7:00 a.m. Sunday, Oct. 25, 1998 to 7:00 a.m. Monday, Oct. 26, 1998.

WEATHER: Sunny and Mild. The temperature range was 33oF to 69oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.2 hours and 883 tons of contaminated soils were thermally treated during the time frame of this report. The majority of down time was attributed to the failure and replacement of the dual feed screw drive chain and the calibration of the weigh belt. The TDF processed contaminated soil at an average rate of 39.8 tons per hour while on line. Related ongoing site activities include feed preparation, debris separation, and rock and ash backfill.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, rock cleaning, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM). Contractor sampled and overpacked (2) drums which were encountered during excavation activities last week. Contractor began collecting scrubber water samples every 4 hours for a 24 hour period. The samples will be analyzed for pH and total suspended solids.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination, and the shipment of used personal protective equipment and miscellaneous site debris to a licensed facility.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for David Modricker (USACE)


Sunday, October 25, 1998

FOR DATES: From 7:00 a.m. Saturday, Oct. 24, 1998 to 7:00 a.m. Sunday, Oct. 25, 1998.

WEATHER: Mostly sunny, mild and breezy. The temperature range was 33oF to 68oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 24.93 hours and 995.8 tons of contaminated soils were thermally treated during the time frame of this report. (Note: A change from daylight savings time occurred during this report period, accounting for one extra hour of operation). The TDF processed contaminated soil at an average rate of 39.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, ash handling, and ash backfilling. Continued with real time perimeter air monitoring.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, October 24, 1998

FOR DATES: From 7:00 a.m. Friday, Oct. 23, 1998 to 7:00 a.m. Saturday, Oct. 24, 1998.

WEATHER: Partly cloudy and cool. The temperature range was 31oF to 62oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.92 hours and 919.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 38.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation of site soil continued on the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) and completed the 24-hour air sampling for BNA at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Gregory A. Beauseigneur for David Modricker (USACE)


Friday, October 23, 1998

FOR DATES: From 7:00 a.m. Thursday, Oct. 22, 1998 to 7:00 a.m. Friday, Oct. 23, 1998.

WEATHER: Overcast and cool. The temperature range was 32oF to 48oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 934.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.4 tons per hour while on line.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant (WWTP) operations, feed preparation, debris handling, debris separation, scrap metal washing, rock cleaning, rock backfilling, ash handling, ash backfilling, and performed a survey of the excavation bottom and the 2-foot rock bottom. A portion of the sheetpiling along the utility corridor was cut down as part of the backfilling process. Excavation continued on the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program and completed the modified OSHA method 93 testing for BNA at the on-site and off-site air monitoring stations.

Received bottom ash and fly ash analytical reports (bot-ash-10-17-98-252 and fly-ash-10-17-98-047) for bin 1 and bin A respectively. Bottom ash test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Fly ash test results indicated that the fly ash materials met project requirements for restricted backfilling (i.e. above elevation 553) due to the concentrations of cadmium, chromium, and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, October 22, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 21, 1998 to 7:00 a.m. Thursday, Oct. 22, 1998.

WEATHER: Cool and overcast with periods of rain. The temperature range was 36oF to 52oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 925.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 38.7 tons per hour while on line.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant (WWTP) operations, feed preparation, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, ash backfilling, and performed a survey of the excavation bottom. Mass excavation continued on the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program and began the modified OSHA method 93 testing for BNA at the on-site and off-site air monitoring stations. Received bottom ash analytical report (bot-ash-10-17-98-252) for bin 2. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, October 21, 1998

FOR DATES: From 7:00 a.m. Tuesday, Oct. 20, 1998 to 7:00 a.m. Wednesday, Oct. 21, 1998.

WEATHER: Clear and cool. The temperature range was 41oF to 62oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 955.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line.

II. ACTIONS TAKEN: The contractor continued TDF operations, waste water treatment plant (WWTP) operations, feed preparation, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. Excavation continued on the southern end of the exclusion zone; two drums were encountered and sampled during excavation activities. Received bottom ash analytical report (bot-ash-10-16-98-251) for bin 3. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547 feet).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs, Larry Newcomer, and Mike Welch (PADEP), a site tour was given to a group of students from Bucknell University.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, October 20, 1998

FOR DATES: From 7:00 a.m. Monday, Oct. 19, 1998 to 7:00 a.m. Tuesday, October 20, 1998.

WEATHER: Clear and mild. The temperature range was 53oF to 77oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 942.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.8 tons per hour while on line.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, waste water treatment plant (WWTP) operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Kenneth R. Hilton for David Modricker (USACE)


Monday, October 19, 1998

FOR DATES: From 7:00 a.m. Sunday, Oct. 19, 1998 to 7:00 a.m. Monday, October 19, 1998.

WEATHER: Clear and mild. The temperature range was 54oF to 78oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 966 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.4 tons per hour while on line.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, waste water treatment plant (WWTP) operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: John Ricard for David Modricker (USACE)


Sunday, October 18, 1998

FOR DATES: From 7:00 a.m. Saturday, Oct. 17, 1998 to 7:00 a.m. Sunday, Oct. 18, 1998.

WEATHER: Mostly sunny. The temperature range was 40oF to 66oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 963.3 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, October 17, 1998

FOR DATES: From 7:00 a.m. Friday, Oct. 16, 1998 to 7:00 a.m. Saturday, Oct. 17, 1998.

WEATHER: Partly cloudy and cool. The temperature range was 39oF to 68oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 942.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 39.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation of site soil continued on the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) and completed the 24-hour air sampling for BNA at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, October 16, 1998

FOR DATES: From 7:00 a.m. Thursday, Oct. 15, 1998 to 7:00 a.m. Friday, Oct. 16, 1998.

WEATHER: Partly cloudy and mild. The temperature range was 42oF to 58oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 900.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 37.7 tons per hour while on line. The contractor continues to screen feed soils and perform debris separation activities.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, ash handling, ash backfilling, rock cleaning, rock backfilling, and operation of the waste water treatment plant (WWTP). Continued with the real time perimeter air monitoring (PAM) program. Conducted 24 hour OSHA 93 and NIOSH 5518 air sampling event at perimeter and offsite stations, as part of the ongoing BNA monitoring program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Continue to evaluate impact of not mixing lime with site soils prior to screening and incineration operations.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Matt Radek, IT/OHM and Sandy Downs, PADEP

PREPARED BY: Chad Thompson for David Modricker (USACE)


Thursday, October 15, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 14, 1998 to 7:00 a.m. Thursday, Oct. 15, 1998.

WEATHER: Partly cloudy and cool. The temperature range was 44oF to 56oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 669.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 28.0 tons per hour while on line. The contractor continues to screen feed soils and perform debris separation activities.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, ash handling, ash backfilling, rock cleaning, rock backfilling, and operation of the waste water treatment plant (WWTP). Excavation continued on the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Mixing lime with contaminated soil prior to incineration will be temporarily discontinued to analyze the impact on incineration feed rates and material handling capabilities.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, October 14, 1998

FOR DATES: From 7:00 a.m. Tuesday, Oct. 13, 1998 to 7:00 a.m. Wednesday, Oct. 14, 1998.

WEATHER: Overcast with evening rain. The temperature range was 46oF to 66oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.6 hours and 635.3 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time to is attributed to feed stoppage as a precaution to power loss during local lightning storm events. The TDF processed contaminated soil at an average rate of 28.1 tons per hour while on line. The low feed rate (less than the target value) continues to be a result of feed soil materials binding to the dual feed screws. USACE and OHM continue to evaluate options to improve the feed rate. Related on-going site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, ash backfilling, drum handling, and operation of the waste water treatment plant. Excavation continued on the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (BOT-ASH-10-09-98-244) for bin 4. The analytical data indicates that the ash is suitable for unrestricted backfilling on site (i.e.: above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, waste water treatment plant (WWTP) operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs, Mike Welch, and Steve Webster (PADEP)

PREPARED BY: David Modricker (USACE)


Tuesday, October 13, 1998

FOR DATES: From 7:00 a.m. Monday, Oct. 12, 1998 to 7:00 a.m. Tuesday, October 13, 1998.

WEATHER: Clear. The temperature range was 45oF to 65oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 620.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 26 tons per hour while on line. The low feed rate (less than the target value) continues to be a result of feed soil materials binding to the dual feed screws. USACE and OHM continue to evaluate options to improve the feed rate.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling, and operation of the waste water treatment plant. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report number BOT-ASH-10-08-98-243. The analytical data indicates that the ash is suitable for unrestricted backfilling on site (i.e.: above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, waste water treatment plant (WWTP) operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: David Modricker (USACE)


Monday, October 12, 1998

FOR DATES: From 7:00 a.m. Sunday, Oct. 11, 1998 to 7:00 a.m. Monday, October 12, 1998.

WEATHER: Clear. The temperature range was 45.3oF to 73.6oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 648.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.1 tons per hour while on line. The low feed rate (less than the target value) continues to be a result of feed soil materials binding to the dual feed screws. USACE and OHM continue to evaluate options to improve the feed rate.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. Received no bottom ash or fly ash analytical reports for this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, waste water treatment plant (WWTP) operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: John Ricard for David Modricker (USACE)


Sunday, October 11, 1998

FOR DATES: From 7:00 a.m. Saturday, Oct. 10, 1998 to 7:00 a.m. Sunday, Oct. 11, 1998.

WEATHER: Overcast with rain in the morning, clearing later in the day. The temperature range was 48.2oF to 61.7oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 824.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.6 tons per hour while on line. USACE and OHM continue to investigate options to improve the feed rate. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. The contractor recovered the weekly OSHA method 93 air samples for BNA at the PAM and offsite ambient air monitoring stations. Received no bottom ash or fly ash analytical reports for this reports coverage.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, October 10, 1998

FOR DATES: From 7:00 a.m. Friday, Oct. 9, 1998 to 7:00 a.m. Saturday, Oct. 10, 1998.

WEATHER: Overcast with rain. The temperature range was 53oF to 56oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 24 hours and 649.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.1 tons per hour while on line. OHM has indicated that the low feed rate (less than the target value) is a result of feed soil materials binding to the dual feed screws. The USACE and OHM are currently investigating the physical properties of the feed soil in an effort to improve the feed rate. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program. Began the weekly OSHA method 93 air sampling event for BNA at the PAM and offsite ambient air monitoring stations. Received bottom ash analytical reports (bot-ash-10-3-98-240, bot-ash-10-4-98-241, bot-ash-10-5-98-242) for bins 1, 7, and 6 respectively. The test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, October 9, 1998

FOR DATES: From 7:00 a.m. Thursday, Oct. 8, 1998 to 7:00 a.m. Friday, Oct. 9, 1998. WEATHER: Rainy and overcast. The temperature range was 53oF to 65oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). Hazardous waste feed to the incinerator resumed at 0257 hours on Friday, October 9, 1998 once the TDF maintenance and calibrations had been completed. The TDF was on-line for 3.4 hours and 112.6 tons of contaminated soils were processed during the time frame of this report. The resumption of feed was delayed due to a local power interruption in the early evening, approximately 1855 hours. The interruption of power to the TDF caused the Thermal Relief Valve (TRV) to open. Power was quickly restored to the site through the TDF's emergency generator and the TRV was closed. The TRV opened again at 2156 hours when power was switched back to the local electrical grid. For both TRV openings, feed to the incinerator had not yet been resumed. The contractor notified the necessary authorities about both TRV openings. Once feed had resumed, additional down time to the incinerator was primarily attributed to the performance of additional maintenance to the feed screw drive. Related on-going site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF preventative maintenance. Once all of the maintenance tasks had been completed, the contractor performed the monthly interlock checks to test the TDF's automatic shutdown systems. The contractor also continued WWTP operations, feed preparation, lime mixing, debris handling, debris separation, ash handling and ash backfilling. Due to steady rain, no new excavation was performed on the southern face of the exclusion zone. The contractor continued with the real time perimeter air monitoring (PAM) program. The fly ash analytical report (fly-ash-10-02-98-045) for bin 'C' was received. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling (i.e. above elevation 553) on-site due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for Frank Cleaver (USACE)


Thursday, October 8, 1998

FOR DATES: From 7:00 a.m. Wednesday, Oct. 7, 1998 to 7:00 a.m. Thursday, Oct. 8, 1998.

WEATHER: Overcast with evening rain. The temperature range was 59oF to 63oF.

I. CURRENT SITUATION: Thermal Destruction Facility (TDF) is currently heating up to specified temperature in order to resume hazardous waste feed to the incinerator later on this date. Once the incinerator reaches the operating temperature, the contractor plans to perform a TDF monthly interlock check in order to test the automatic shutdown systems of the incinerator. The contractor continues to work 24 hours a day, seven days a week performing maintenance on the TDF equipment. While the incinerator is off-line, the contractor continues to excavate, screen debris, and prepare and stockpile feed for the resumption of operations.

II. ACTIONS TAKEN: The contractor continued TDF preventative maintenance. Maintenance tasks that were completed included the installation of the new feed screw assembly, the replacement of the bent flights on the Dry Ash Drag, the replacement of the expansion joints of the acid gas scrubber system's piping, and several other preventative maintenance tasks. The contractor also continued WWTP operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling and ash backfilling. Excavation near former sediment basin 1 on the western side of the exclusion zone resumed on this date. Rusmar foam, an odor suppressant, was sprayed on the excavation face upon completion of the day's work. MRI, the subcontracted air monitoring firm, continued with the real time perimeter air monitoring (PAM) program and finished the monthly TSP/metals sampling at each of the PAM stations on-site. Received bottom ash analytical reports for (bot-ash-09-30-98-237, bot-ash-10-01-98-238, bot-ash-10-02-98-239) for bins 4, 3, and 2 respectively. Analytical test results indicated that the thermally treated soils met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Resumption of feed to the incinerator is tentatively scheduled to resume on Thursday, October 8, 1998 in the early evening.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: A group of Junior Engineers in Training for the USACE toured the site.

PREPARED BY: Steve Gibson for Bill Werntges (USACE)


Wednesday, October 7, 1998

FOR DATES: From 7:00 a.m. Tuesday, Oct. 6, 1998 to 7:00 a.m. Wednesday, Oct. 7, 1998.

WEATHER: Overcast and cool. The temperature range was 44oF to 61oF.

I. CURRENT SITUATION: Thermal Destruction Facility (TDF) remained off-line in a cold standby for the period of this report to allow for the replacement of the kiln feed screws and their external components. The contractor is working 24 hours a day, seven days a week performing maintenance to the TDF equipment including replacement of covers on the dry ash drag and expansion joints in the acid gas scrubber piping. While the incinerator is off-line, the contractor continues to excavate, screen debris, and prepare and stockpile feed for the resumption of operations.

II. ACTIONS TAKEN: The contractor continued TDF preventative maintenance, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling and ash backfilling. Excavation near former sediment basin 1 on the southern end of the exclusion zone resumed on this date. Rusmar foam, an odor suppressant, was sprayed on the excavation face upon completion of the day's work. MRI, the subcontracted air monitoring firm, continued with the real time perimeter air monitoring (PAM) program and began the monthly TSP/metals sampling at each of the PAM stations on-site.

There was a planned Thermal Relief Valve (TRV) opening at 1000 hours. The TRV was opened in order to allow workers to perform maintenance on the acid gas scrubber tanks and piping. Feed to the incinerator had been off for over 11 hours and the proper authorities were notified by the contractor prior to the TRV opening.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

Resumption of feed to the incinerator is tentatively scheduled to resume on Friday, October 9, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Mike Welch (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, October 6, 1998

FOR DATES: From 7:00 a.m. Monday, Oct. 5, 1998 to 7:00 a.m. Tuesday, Oct. 6, 1998.

WEATHER: Overcast and clearing in the evening. The temperature range was 43oF to 62oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week performing maintenance to the Thermal Destruction Facility (TDF). The TDF is currently in warm shut down mode in order to replace the feed screws and perform other preventative maintenance tasks. Prior to the shutdown, the TDF was on-line for 15.6 hours and 512.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 32.8 tons per hour while on line.

Feed was stopped at approximately 2257 hours on October 5, 1998, to replace a broken expansion joint on a 2-inch water line on the discharge side of the acid gas scrubber pump. Feed has not resumed. See notes under section IV below.

II. ACTIONS TAKEN: The contractor continued TDF operations and maintenance, WWTP operations, feed preparation, lime mixing, debris handling, debris separation, ash handling and ash backfilling. No new excavation was performed on the southern end of the exclusion zone. The used carbon from the feed preparation building air handling unit was shipped off-site to a carbon regeneration plant. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

There will be a planned Thermal Relief Valve (TRV) opening on Tuesday, October 6, 1998 at approximately 1000 hours. The TRV will be opened in order to perform preventative maintenance on the internal components of the TDF. Feed to the incinerator is tentatively planned to restart October 8, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: An expansion joint broke on a water pipe that is part of the system that maintains the water levels in the acid gas scrubber. As a result, the contractor initiated corrective action that included an unplanned TRV opening in order to prevent the acid gas scrubber tanks from overheating. The TRV was opened at 2314 hours. Feed to the incinerator had been shut down approximately 17 minutes prior to the opening. OHM made the necessary notifications as a result of the TRV opening. The TRV was closed at 0014 hours on Tuesday morning. Feed to the incinerator was not restarted, and the maintenance, which was planned to begin later in the morning, was initiated.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, October 5, 1998

FOR DATES: From 7:00 a.m. Sunday, Oct. 4, 1998 to 7:00 a.m. Monday, Oct. 5, 1998.

WEATHER: Overcast, cool, rain. The temperature range was 47oF to 50oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.2 hours and 785.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.8 tons per hour while on line. Downtime was attributable to the feed screws tripping due to soil sticking to them and the replacement of a fitting on the line that measures differential pressure on the exit side of the baghouse. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, and ash handling. Collected OSHA93 and NIOSH 5518 sampling pumps at PAM and OAM stations as part of the on going BNA air monitoring program. Continued with the real time perimeter air monitoring (PAM).

No new ash analytical reports received during the period covered by this report. Due to extremely wet conditions in the field, the backfill of ash was curtailed and will be resumed as conditions allow.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. A vacuum truck is on-site Monday October 5, 1998 to pick up the used granular carbon from the feed prep air handling unit. The used carbon will be sent off-site for regeneration. There is a planned shut down of the TDF scheduled for October 6th and 7th . During the planned shut down the contractor plans to perform several preventative maintenance tasks which include the replacement of the feed screws. During this maintenance period, a thermal relief vent (TRV) opening is planned.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for David Modricker (USACE)


Sunday, October 4, 1998

FOR DATES: From 7:00 a.m. Saturday, Oct. 3, 1998 to 7:00 a.m. Sunday, Oct. 4, 1998.

WEATHER: Overcast, cool, rain. The temperature range was 43 oF to 51oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 818.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM).

Received bottom ash analytical report (bot-ash-09-29-98-236) for bin 5. Test results indicated that the thermally treated soil met the project requirements for restricted backfilling (i.e. above elevation 553) due to the level of arsenic being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a planned shut down of the TDF scheduled for October 6th and 7th of next week. During the planned shut down the contractor plans to perform several preventative maintenance tasks which include the replacement of the feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for David Modricker (USACE)


Saturday, October 3, 1998

FOR DATES: From 7:00 a.m. Friday, Oct. 2, 1998 to 7:00 a.m. Saturday, Oct. 3, 1998.

WEATHER: Clear and cool. The temperature range was 38oF to 61oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 827.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM).

Received bottom ash analytical report (bot-ash-09-28-98-235) for bin 6. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-09-27-98-044) for bin 'B'. Test results indicated that the fly ash materials met the project requirements for restricted backfilling (i.e. above elevation 553) due to the levels of selenium, chromium, and lead being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a planned shut down of the TDF scheduled for October 6th and 7th of next week. During the planned shut down the contractor plans to perform several preventative maintenance tasks which include the replacement of the feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, October 2, 1998

FOR DATES: From 7:00 a.m. Thursday, October 1, 1998 to 7:00 a.m. Friday, October 2, 1998.

WEATHER: Breezy and cool. The temperature range was 38oF to 63oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 828.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.7 tons per hour while on line.

The contractor has indicated that the reduced hourly feed rate (less than the target) is a result of feed material adhering to the feed screws. USACE and OHM are investigating soil feed characteristics and the feed screw configuration in an effort to optimize the feed rate. Although the feed rate while on line has been below the target since the middle of September, the overall productivity in September was above the target.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, stockpile grading, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed at the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-09-27-98-234) for bin 7. Test results indicated that the thermally treated soil met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation of contaminated material, drum handling (if drums are encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The TDF is currently scheduled to be off line on October 6th and 7th in order to perform several maintenance tasks, including the replacement of the kiln feed screws. A planned thermal relief valve (TRV) opening is scheduled during the maintenance activities next week.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Thursday, October 1, 1998

FOR DATES: From 7:00 a.m. Wednesday, September 30, 1998 to 7:00 a.m. Thursday, October 1, 1998.

WEATHER: Sunny and mild. The temperature range was 53oF to 80oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 827.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.6 tons per hour while on line. The contractor has indicated that the reduced feed rate (less than the target) is a result of feed material adhering to the feed screws. USACE and OHM are investigating soil feed characteristics and the feed screw configuration in an effort to optimize the feed rate. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, lime mixing, stockpile grading, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. No new excavation was performed at the southern end of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-09-26-98-233) for bin 1. Test results indicated that the thermally treated soil met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. The TDF is currently scheduled to be off line sometime next week (tentatively October 6th and 7th ) to perform several maintenance tasks, including the replacement of the kiln feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Raymond Pompe, Pete Ives, Matt Radek (OHM-IT)

PREPARED BY: Steve Gibson for Dave Modricker (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.